OpenStudy (anonymous):

sqrt2 sin^2-sin=0 solve the equation

OpenStudy (anonymous):

You mean: \[\sqrt{2\sin^2(x) - \sin(x)} = 0\]

OpenStudy (anonymous):

@ashleedean12 want to reply or not??

OpenStudy (anonymous):

\[\sqrt{2}\sin ^{2}\theta-\sin \theta=0\]

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Where were you??

OpenStudy (anonymous):

working on a practice test

OpenStudy (anonymous):

What you can do here, just take one sin(theta) out of the equation, I mean to factor out one sin(theta).. Can you do that??

OpenStudy (anonymous):

i have no idea how to do that

OpenStudy (anonymous):

I am just giving you an example: \[\cos^2(\theta) - \cos(\theta) = 2 \implies \color{red}{\cos(\theta) (\cos(\theta) - 1 ) = 2}\]

OpenStudy (anonymous):

Like this taking one common out of the equation.. Same you can apply there in your question.

OpenStudy (anonymous):

@yummydum will explain you better than me...

OpenStudy (anonymous):

i havent done this in a while but i think this is how: \[\sqrt{2}\sin^2\theta-\sin\theta=0\]\[\sin\theta(\sqrt{2}\sin\theta-1)=0\]\[\sin\theta=0~~~~~~~~~~~~~~~~~~~~\sqrt{2}\sin\theta-1=0\]\[\theta=\sin^{-1}(0)~~~~~~~~~~~~~~\sqrt{2}\sin\theta=1\]\[\theta=0,\pi+\pi n~~~~~~~~~~~~~~\sin\theta={1\over\sqrt{2}}\]\[~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\theta=\sin^{-1}\left(1\over\sqrt{2}\right)\]\[~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\theta={\pi\over4}+2\pi n, {3\pi\over4}+\pi n\] and so, \(\theta=0,~\pi+\pi n,~{\pi\over4}+2\pi n,~{3\pi\over4}+\pi n\)

OpenStudy (anonymous):

so sqrt2 -1sin -sin=0

OpenStudy (anonymous):

no, look at what i did

OpenStudy (anonymous):

that confuses me

OpenStudy (anonymous):

so would the answer be 0,pi, pi/4, 3pi/4

OpenStudy (anonymous):

Ha ha ha ha... @yummydum

OpenStudy (anonymous):

Don't go with the answers @ashleedean12 Just understand the solution that @yummydum has given above..

OpenStudy (anonymous):

you factor out a \(\sin\theta\) and now you have two factors, \(\sin\theta\) and \(\sqrt{2}\sin\theta-1\), both of which are equal to \(0\) separately set the factors equal to \(0\) and solve.....you have to know the unit circle to know what \(\theta\) is at \(0\) and \({1\over\sqrt{2}}\)

OpenStudy (anonymous):

\[\large \color{red} {and~~so,~~\theta=0,~\pi+\pi n,~{\pi\over4}+2\pi n,~{3\pi\over4}+\pi n}\]